OpenStudy (anonymous):

2=log x +log (x+48) solve

OpenStudy (anonymous):

ok, so log(a) + log(b) = log(ab) this means we can group together on the RHS 2 = log(x(x+48)) now we want to get rid of the logs assuming this is the common base 10 logarithm we raise 10 to the power of each side: 10^2 = 10^log(x(x+48)) now the log and the 10^ cancel leaving 100 = x(x+48) do you follow the steps up to there?

OpenStudy (anonymous):

can't be more clear than that...

OpenStudy (anonymous):

:D i could have explained why they cancel i suppose

OpenStudy (anonymous):

Thanks Still a little confused

OpenStudy (precal):